# ToonifyGB: StyleGAN-based Gaussian Blendshapes for 3D Stylized Head Avatars

### v1

@@ 4.3. Implementation Details
Figure 4. Visualization of stylized video generation: We present details of the real head from the input video, and the “Arcane” stylized head generated by our method. From left to right, the results for the data samples “bala” and “wojtek_1” are shown. The generated results are highly reproducible, preserving key features of the original head, such as the beard, mouth shape, and the black mole above the eye.
[…]
First row focuses on sides, second row on details.
@@
Table 1. Duration (in seconds) of the input videos, and inference time (in seconds) of our method. The input videos are $512\times 512$ pixels. The inference is done on an RTX 4090 GPU.
@@
Table 2. Quantitative comparison of Video Stabilization: We compare the original input (OI), the aligned input (AI), our “Arcane” (OA), and the aligned “Arcane” (AA) videos.
@@
4.4. Quantitative Comparison
4.4.1. Video Stabilization
We employ an improved StyleGAN model to generate six videos in the “Arcane” style. The duration of each video and the corresponding inference time are presented in Table 1. For videos lasting between 60 and 160 seconds, the generation inference time ranges from approximately 100 to 350 seconds.
We adopt a common facial alignment method using the facial keypoint predictor (Kazemi and Sullivan, 2014) for input videos. We compare the original input videos (OI) with the aligned input videos (AI), and compare the “Arcane” style videos generated by our method (OA) with those generated from AI (AA). As shown in Table 2, both IFT and ISI values of AI are consistently lower than those of OI. Similarly, AA obtain lower IFT and ISI values than OA. These results indicate that applying facial alignment and cropping before frame-by-frame generation (i.e., AI and AA) produces more unstable (i.e., more jittery) videos.
Table 3. Quantitative comparison of 3D head avatars: We evaluate our method and state‑of‑the‑art methods on the INSTA (Zielonka et al., 2023) dataset. In each metric group, the best value is highlighted in bold, and the second‑best is underlined.
@@ 4.4.1. Video Stabilization
Table 4. Performance comparison: We record the training time (in minutes) and the rendering speed (in fps) of 3DGB and our method in both “Arcane” (A) and “Pixar” (P) styles.
@@
4.4.2. 3D Head Avatar
@@
In addition, although our method integrates stylization into 3D head avatars, its training and rendering times remain comparable to those of the method of 3DGB. In certain cases, our method is even more efficient in both training and rendering. Combined with the additional time required for video generation (as shown in Table 1), the overall time cost of our method remains acceptable.
Figure 5. Qualitative comparison of each stage: We present the input video head frames, the corresponding stylized videos, and the 3D head avatars synthesized by our method.
@@
Qualitative comparison between video and 3D head avatars.
Figure 6. Qualitative comparison of 3D Gaussian head avatars: We present 3D head avatars using Gaussian blendshapes synthesized by 3DGB (Ma et al., 2024) and our method.
[…]
Qualitative comparison between 3DGB and ours.
4.5. Qualitative Comparison
We present the original video head frames, their corresponding stylized versions generated by our method, and the 3D stylized head avatars synthesized using Gaussian blendshapes. The qualitative comparison is shown in Figure 5. The examples are selected from the “bala” data in the “Pixar” style and the “wojtek_1” data in the “Arcane” style. In the stylized video, the “bala” data exhibits artifacts along the edge of the side of the head. We attribute this to the latent space distribution learned by StyleGAN, which tends to produce striped artifacts when the viewing angle falls outside the distribution covered by the training data. Notably, such artifacts are not present in the corresponding 3D stylized head avatars during rendering. Furthermore, the 3D stylized head avatars successfully preserve fine details from the stylized videos, such as the mole near the eye in the “wojtek_1” data. However, since the 3D avatar synthesis mainly focuses on the facial region, the neck area is typically blurred, as observed in both examples. This blurring also leads to the lower quantitative performance, as the neck region is included in the evaluation.
The qualitative comparison with 3DGB (Ma et al., 2024) is presented in Figures 6 and 11. Our method effectively captures and preserves high-frequency details in the stylized videos. Compared to the method of 3DGB, ToonifyGB can synthesize 3D stylized head avatars with comparable quality and detail.
4.6. Visualization
4.6.1. Generated Video Details
@@
4.6.2. Gaussian Blendshapes
We present eight Gaussian blendshapes for the “justin” sample, as well as their corresponding mesh blendshapes, as shown in Figure 10. For more visualizations, please refer to the video in our supplementary material.
Table 5. Ablation Study: We compare 3D head avatars (using “justin” data generated in both styles) synthesized from different input videos: one generated by our method, and the other using facial alignment and cropping as prprocessing.
[…]
w/ Face Align & Crop 32.23 0.9387 0.1587
Ours 33.27 0.9645 0.0796
Figure 7. Ablation Study: We present the 3D head avatar synthesized by our method, as well as that synthesized using the cross-identity reenactment from 3DGB (Ma et al., 2024).
[…]
Ablation study on our method and cross-identity reenactment from 3DGB.
4.7. Ablation Study
4.7.1. Facial Alignment and Cropping
@@
We compare 3D stylized head avatars synthesized from videos stylized by our pipeline with those synthesized from videos preprocessed by facial alignment and cropping. We evaluate the resulting avatars using PSNR, SSIM, and Learned Perceptual Image Patch Similarity (LPIPS) metrics. The results are shown in Table 5. The avatars synthesized by our method outperform those synthesized from videos with facial alignment and cropping as preprocessing in all evaluation metrics.
4.7.2. Cross-identity Reenactment
One of the key contributions of this work is the real-time synthesis of stylized head avatar animations using expression blendshapes for each frame. The method of 3DGB (Ma et al., 2024) supports cross-identity reenactment, enabling the transfer of facial expressions between different head avatars while preserving the target’s personal attributes. However, transferring the expressions from a real head to a rendered stylized head avatar, as shown in Figure 7, may produce unsatisfactory results, particularly around the mouth region. This error occurs because the source heads are stylized, and the expression blendshapes differ significantly. This demonstrates the importance of video-based acquisition, as 3DGB method based on cross-identity reenactment is insufficient for efficient animation synthesis.
@@ 4.7.2. Cross-identity Reenactment
Figure 8. Limitation: We present side-view renderings synthesized by 3DGB and our method.
[…]
Side View of 3DGB and ours on “justin” data.
5. Limitation and Future Work
@@
Our method struggles to render side views of 3D stylized head avatars if the training data (i.e., input video) lacks side views of the real head. As shown in Figure 8, this limitation also affects the baseline method. In fact, existing NeRF-based and Gaussian-based methods have yet to effectively address this problem. Rendering novel views from single view training data remains an open problem for future work. Employing 2D GANs to synthesize videos with side views as additional training data and enhancing the generalization ability of our model are two potential solutions, which are interesting directions for future investigation.
6. Conclusion
@@
We propose a novel two-stage framework, ToonifyGB, which uses Gaussian blendshapes to synthesize the head animations in diverse toonified styles from input videos. We incorporate an improved StyleGAN model to generate stylized videos beyond facial alignment and cropping as preprocessing in Stage 1. This process produces more stable video outputs, providing a basis for high-quality 3D head avatar animation synthesis. ToonifyGB serves as a bridge between 2D generation and 3D modeling, offering valuable insights into applying 2D GANs for 3D head reconstruction.

### v2

@@ Introduction
Figure 3: Visualization of stylized video generation: We present details of the real head from the input video, and the “Arcane” stylized head generated by our method. From left to right, the results for the data samples “bala” and “wojtek_1” are shown.
Implementation Details
To ensure a fair performance comparison, the training and testing of all methods are performed on a single RTX 4090 GPU. Our methods are implemented in Python using the PyTorch framework. For 2D stylized video generation, we use the pre-trained models provided by StyleGANEX (Yang et al. 2023). For training the 3D stylized head avatars, we employ the Adam optimizer (Kingma and Ba 2015), setting the initial learning rates of the Gaussian properties $\{\mathbf{x}_{k},\alpha_{k},\mathbf{s}_{k},\mathbf{q}_{k},SH_{k}\}$ to $3.2\times 10^{-7}$, $5\times 10^{-5}$, $5\times 10^{-4}$, $1\times 10^{-4}$, and $1.25\times 10^{-3}$, respectively. Following 3DGB (Ma et al. 2024), the initial number of sampled Gaussians is 50k for the neutral head model and 14k for the mouth interior.
@@
Table 1: Duration (in seconds) of the input videos, and inference time (in seconds) of our method.
@@
Table 2: Quantitative comparison of video stabilization: We compare the original input (OI), the aligned input (AI), our “Arcane” (OA), and the aligned “Arcane” (AA) videos.
@@
Table 3: Quantitative comparison of 3D head avatars: We evaluate our method and SOTA methods on the INSTA (Zielonka, Bolkart, and Thies 2023) dataset. In each metric group, the best value is highlighted in bold, and the second‑best is underlined.
@@
Table 4: Performance comparison: We record the training time (in minutes) and the rendering speed (in fps) of 3DGB and our method in both “Arcane” (A) and “Pixar” (P) styles.
@@
Quantitative Comparison
Video Stabilization
We adopt an improved StyleGAN model to generate six videos in the “Arcane” style. The durations of the videos and their corresponding inference times are summarized in Table 1. All input videos have a resolution of $512\times 512$ pixels, and inference is performed on a single NVIDIA RTX 4090 GPU. For video durations ranging from 60 to 160 seconds, the generation times span approximately 100 to 350 seconds.
To evaluate the impact of preprocessing, we apply a standard face alignment technique based on a facial keypoint predictor (Kazemi and Sullivan 2014) to the input videos. We compare the original input videos (Original Input, OI) with their aligned counterparts (Aligned Input, AI). Likewise, we compare the “Arcane” style outputs generated from unaligned inputs (Ours Arcane, OA) with those generated from aligned inputs (Aligned Arcane, AA). As shown in Table 2, both the Inter-frame Transformation Fidelity (ITF) and Inter-frame Similarity Index (ISI) scores for AI are consistently lower than those for OI. Similarly, AA exhibits lower ITF and ISI scores compared to OA. These results suggest that applying face alignment and cropping prior to frame-by-frame generation (i.e., AI and AA) tends to introduce greater temporal instability, resulting in more jittery outputs.
3D Head Avatar
@@
In addition, although our method integrates stylization into 3D head avatars, its training and rendering times remain comparable to those of the method of 3DGB (Ma et al. 2024). In certain cases, our method is even more efficient in both training and rendering. Combined with the additional time required for video generation (as shown in Table 1), the overall time cost of our method remains acceptable.
Figure 4: Qualitative comparison of each stage: We present the input video head frames, the corresponding stylized videos, and 3D head avatars synthesized by our method.
Figure 5: Qualitative comparison of baseline and ours: We present 3D head avatars using Gaussian blendshapes synthesized by 3DGB (Ma et al. 2024) and our method.
Qualitative Comparison
We present the original video head frames, the corresponding stylized video frames generated by our method, and the 3D stylized head avatars synthesized using Gaussian blendshapes. The qualitative comparison is shown in Figure 4. The examples are selected from the “bala” dataset in the “Pixar” style and the “wojtek_1” dataset in the “Arcane” style. In the stylized video, the “bala” data exhibits artifacts along the side edge of the head. We attribute this to the latent space distribution learned by StyleGAN, which tends to produce striped artifacts when the viewing angle falls outside the distribution covered by the training data. Notably, these artifacts are not present in the corresponding 3D stylized head avatars rendered by our method. Furthermore, the 3D stylized head avatars successfully preserve fine details from the stylized videos, such as the mole near the eye in the “wojtek_1” dataset. However, since the 3D avatar synthesis mainly focuses on the facial region, the neck area is typically blurred, as observed in both cases. This blurring leads to the lower quantitative performance, since the neck region is included in the evaluation.
The qualitative comparison with 3DGB (Ma et al. 2024) is presented in Figures 5. Our method effectively captures and preserves high-frequency details in the stylized videos. Compared to the SOTA method, ToonifyGB can synthesize 3D stylized head avatars with comparable quality and detail.
Table 5: Ablation study on face alignment and cropping: We compare 3D head avatars synthesized from different input videos: one generated by our method, and the other using face alignment and cropping as prprocessing.
@@
Face Align & Crop 32.23 0.9387 0.1587
Ours 33.27 0.9645 0.0796
Figure 6: Ablation study on the effect of different driving videos: We present 3D stylized head avatar animation driven by the original input videos and our generated videos.
Visualization
Generated Video Details
@@
Ablation Study
Face Alignment and Cropping
@@
We compare 3D stylized head avatars (using the “justin” data) synthesized from videos processed by our method against those generated from videos preprocessed with face alignment and cropping. The resulting avatars are evaluated using PSNR, SSIM, and Learned Perceptual Image Patch Similarity (LPIPS). As shown in Table 5, our method outperforms the traditional method with face alignment and cropping across all evaluation metrics. This demonstrates that our method effectively eliminates jitter during video generation, enabling higher-quality synthesis of 3D stylized head animations.
Source Videos for Driving Animation
Compared with the architecture of 3DGB (Ma et al. 2024) that synthesizes 3D photo-realistic head avatars, our framework includes an additional Stage 1 to generate the stylized video. To demonstrate the importance of the generated stylized video in driving the animation, we compare the results of using the original input video (real face) versus our generated stylized video as the driving source, as shown in Figure 6.
It can be observed that using the original input video (real face) as the driving source often leads to unsatisfactory results, especially around the mouth region. This error occurs due to significant differences in expression blendshapes between the real and stylized domains. These results highlight the importance of the stylized videos generated by Stage 1 of our framework. Therefore, we recommend using the generated stylized videos, rather than the original input videos, as the driving source for 3D stylized head avatar animation.
@@
Figure 7: Limitation: We present side-view renderings synthesized by 3DGB (Ma et al. 2024) and our method.
Limitation
@@
Our method struggles to render side views of 3D stylized head avatars when the training data (i.e., input video) lacks side-view representations of the real head. As shown in Figure 7, we present side-view renderings synthesized by both 3DGB (Ma et al. 2024) and our method, and this limitation is also observed in the state-of-the-art methods. In fact, existing NeRF-based and Gaussian-based methods have yet to effectively address this issue. Rendering novel views from single-view training data remains an open problem for future research. Two directions to address this limitation include employing 2D GANs to synthesize videos with side views as additional training data, and enhancing the generalization ability of our model.
Conclusion
@@
We propose a novel two-stage framework, ToonifyGB, which utilizes Gaussian blendshapes to synthesize head animations in diverse styles from monocular videos. In Stage 1, the proposed method adopts an improved StyleGAN-based model to generate stylized videos without requiring face alignment or cropping as preprocessing. This results in more temporally stable outputs, providing a reliable foundation for high-quality 3D head avatar animation synthesis. Stage 2 focuses on constructing 3D stylized head avatars using Gaussian blendshapes, enabling fine-grained expression modeling and satisfactory animation. Our method supports real-time generation of stylized avatar animations in popular styles such as “Arcane” and “Pixar”.
For future work, we plan to integrate motion capture technologies to enable real-time expression control of 3D stylized avatars. This direction is expected to further broaden the applicability of ToonifyGB in virtual character interaction and personalized avatar generation.